# Simulating Roman+Gaia Combined Astrometry, Parallaxes, and Proper Motions

### v1

@@ VI Conclusions
We have created a realistic simulation tool that combines Gaia with Roman to predict uncertainties in position, parallax, and PM as a function of magnitude given a user-provided Roman observing strategy (Sections II & III, Figure 3). This tool is actively being used by multiple groups writing Roman proposals to ensure that various astrometric thresholds for scientific goals are being met. Our code is publicly available with additional Jupyter notebooks providing tutorials for its use. In this work, we have simulated the core surveys of the Roman mission (Section V): GPS, HLTDS, GBTDS, and HLWAS. We predict that the HLTDS and GBTDS will provide high quality parallaxes and PMs to much fainter magnitudes than Gaia alone will ever see (Figures 8 to 11). We also find that the astrometric quality of the HLWAS is extremely dependent upon the final selected observing strategy (Section V.4 and Figure 12). We show that choosing a Roman baseline of 6 months for HLWAS leads to almost useless PM uncertainties ( mas/yr) for mag, while collecting observations over 4.5 years results in PMs that are 1000 times more precise. It is our strongest desire that the committees in charge of the Roman timeline considers the results of these simulations when designing the final schedule.
@@
This work was carried out in collaboration with the HSTPROMO (High-resolution Space Telescope PROper MOtion) Collaboration888https://www.stsci.edu/~marel/hstpromo.html, a set of projects aimed at improving our dynamical understanding of stars, clusters, and galaxies in the nearby Universe through measurement and interpretation of proper motions from HST , Gaia , JWST , Roman , and other space observatories. We thank the collaboration members for sharing their ideas and software.
@@ Appendix A Line-of-Sight Velocity Impact on Astrometry
Putting in some typical values – such as km/s, years, mas/yr, and mas – we find that the stellar distance must be much larger than pc to be confident that the LOS velocity will not significantly change the parallax or PM over a 10 year baseline.

### v2

@@ VI Conclusions
We have created a realistic simulation tool that combines Gaia with Roman to predict uncertainties in position, parallax, and PM as a function of magnitude given a user-provided Roman observing strategy (Sections II & III, Figure 3). The statistics at the heart of our tool can easily be generalized to include astrometric positions from other surveys and datasets. This tool is actively being used by multiple groups writing Roman proposals to ensure that various astrometric thresholds for scientific goals are being met. Our code is publicly available with additional Jupyter notebooks providing tutorials for its use. In this work, we have simulated the core surveys of the Roman mission (Section V): GPS, HLTDS, GBTDS, and HLWAS. We predict that the HLTDS and GBTDS will provide high quality parallaxes and PMs to much fainter magnitudes than Gaia alone will ever see (Figures 10 to 13). We also find that the astrometric quality of the HLWAS is extremely dependent upon the final selected observing strategy (Section V.4 and Figure 14). We show that choosing a Roman baseline of 6 months for HLWAS leads to almost useless PM uncertainties ( mas/yr) for mag, while collecting observations over 4.5 years results in PMs that are 1000 times more precise. It is our strongest desire that the committees in charge of the Roman timeline considers the results of these simulations when designing the final schedule.
@@
This work was carried out in collaboration with the HSTPROMO (High-resolution Space Telescope PROper MOtion) Collaboration111111https://www.stsci.edu/~marel/hstpromo.html, a set of projects aimed at improving our dynamical understanding of stars, clusters, and galaxies in the nearby Universe through measurement and interpretation of proper motions from HST , Gaia , JWST , Roman , and other space observatories. We thank the collaboration members for sharing their ideas and software.
@@ Appendix A Line-of-Sight Velocity Impact on Astrometry
Putting in some typical values – such as km/s, years, mas/yr, and mas – we find that the stellar distance must be much larger than pc to be confident that the LOS velocity will not significantly change the parallax or PM over a 10 year baseline. For some of the highest-precision PMs and parallaxes expected from next-generation telescopes, astrometric uncertainties like as/yr and as will be possible (e.g., Figure 12). In these cases, the distance threshold would increase to pc (i.e. a factor of times larger).